# N-Agent Ad Hoc Teamwork

### v2

@@ Abstract
Current approaches to learning cooperative multi-agent behaviors assume relatively restrictive settings. In standard fully cooperative multi-agent reinforcement learning, the learning algorithm controls all agents in the scenario, while in ad hoc teamwork, the learning algorithm usually assumes control over only a single agent in the scenario. However, many cooperative settings in the real world are much less restrictive. For example, in an autonomous driving scenario, a company might train its cars with the same learning algorithm, yet once on the road, these cars must cooperate with cars from another company. Towards expanding the class of scenarios that cooperative learning methods may optimally address, we introduce $N$-agent ad hoc teamwork (NAHT), where a set of autonomous agents must interact and cooperate with dynamically varying numbers and types of teammates. This paper formalizes the problem, and proposes the Policy Optimization with Agent Modelling (POAM) algorithm. POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem, that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors. Empirical evaluation on tasks from the multi-agent particle environment and StarCraft II shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
@@ 3 NAHT Problem Formulation
Without loss of generality, let $C(\theta)=\{\pi^{i}_{\theta}(.|s)\}_{i=1}^{N}$ denote a set of $N$ controlled agent policies parameterized by $\theta$, such that a learning algorithm might optimize for $\theta$. Let the $\bm{\pi}^{(M)}$ indicate a team of $M$ agents and $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling such a team from $U$ and $C$ via the team sampling procedure. The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
@@ 4 The Need for Dedicated NAHT Algorithms
Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$. The team wins if the sum of the chosen bits is exactly $1$. We denote the probability of winning by $P(s=1)$. Suppose the uncontrolled agents follow a policy that independently selects $1$ with probability $p=\frac{1}{M}$.  [^3] Lemma A.1 shows that this is the optimal value of $p$ for a team composed of such agents. In the following, we consider the three agent case, $M=3$, for simplicity.
In the AHT problem setting, a learning algorithm assumes control of only a single agent. Let $p_{\text{aht}}$ denote the probability with which the AHT agent selects 1. Given the aforementioned team of uncontrolled agents, we show that any value of $p_{\text{aht}}$ results in the same $P(s=1)$, i.e. all possible policies for the AHT agent are optimal. This occurs because the probability of winning, $P(s=1)=\frac{4}{9}$, is independent of $p_{\text{aht}}$ (Lemma A.2).
@@
Finally, we show that there exists an NAHT policy that controls both agents and obtains a higher winning probability. Consider the policy where one controlled agent always plays 0, while the other plays 1 with probability $p_{\text{naht}}$. We show that the optimal $p_{\text{naht}}=1$, and the probability of winning $P(s=1)=\frac{2}{3}>\frac{4}{9}$ (Lemma A.4). Thus, we have exhibited an NAHT scenario where an AHT policy that is optimal when $N=1$, performs worse than a simple NAHT joint policy in the $N=2$ setting.
@@ 5 Policy Optimization with Agent Modeling (POAM)
Figure 2: POAM trains a single policy network, which characterizes the behavior of all controlled agents (green), while uncontrolled agents (yellow) are drawn from $U$. Data from both controlled and uncontrolled agents is used to train the critic, $V_{\pi}$ while the actor $\pi(\theta)$ is trained on data from the controlled agents only. The actor and critic are both conditioned on a learned team encoding vector.
This section describes the proposed Policy Optimization with Agent Modeling (POAM) method to trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates. POAM relies on an agent modeling network to initially build an encoding vector of the collection of teammates encountered during an interaction. Adaptive agent policies that can maximize the controlled agents’ returns are then learned by training a policy conditioned on the environment observation and team encoding vector. The training processes for agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
@@ 5.1 Agent Modeling Network
Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior. However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team. To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team encoding vectors that characterize the observed behavior of teammates in the encountered team.
@@
We identify two main criteria for desirable team encoding vectors. First, team encoding vectors should identify information regarding the unknown state and behavior of other agents in the environment (both controlled and uncontrolled). Second, team encoding vectors should ideally be computable from the sequence of local observations and actions of the team. Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
@@
POAM produces desirable team encoding vectors by training a model with an encoder-decoder architecture, which is illustrated by red components in Figure 2. The encoder, $f^{\text{enc}}_{\theta^{e}}:H^{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team encoding vector, $e^{t}_{i}\in\mathbb{R}^{n}$. This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state that is unavailable under partial observability. The decoder networks, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O^{-i}$ and $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$, are trained to predict the observations and actions of the modeled agents at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the NAHT agents’ decision-making process. Formally, this decoder is trained to minimize the following loss function to accurately predict the modeled agents’ observations and actions:
$L_{\theta^{e},\theta^{o},\theta^{a}}(h^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h^{t}))))\Big{)}.$ (2)
@@ 5.2 Policy and Value Networks
The policy network, $\pi^{i}_{\theta^{p}}:H^{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents. Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents. When training the policy network, we also rely on a value network $V_{\theta^{c}}:H^{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, which measures the expected returns given $h_{t}$, and $e^{t}_{i}$. The value network serves as a baseline to reduce the variance of the gradient updates and also conditions on the learned teammate embeddings, for similar reasons to the policy.
POAM then trains the policy and value networks using an approach based off the Independent PPO algorithm [43] (IPPO). IPPO is selected as the base MARL algorithm for two reasons. First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness. Second, it has been demonstrated to be effective on various MARL tasks. POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
$L_{\theta^{c}}(h^{t})=\dfrac{1}{2}\Big{(}V_{\theta^{c}}(h^{t},f^{\text{enc}}_{\theta^{e}}(h^{t}))-\hat{V}^{t}\Big{)}^{2},$ (3)
where $\hat{V}^{t}=\sum_{l=0}^{T}\gamma^{l}r^{t+l}$. The policy network is trained to minimize the PPO loss function [34].
@@
In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment, where the team consists of a mix of controlled and uncontrolled agents. POAM leverages data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies. This aspect of POAM is a significant departure from PPO, which is a fully on-policy algorithm. However, only data from the controlled agents is used to train the policy network, as the policy update is highly sensitive to off-policy data.
@@ 6 Experiments and Results
  • Q1: Does POAM lead to better sample efficiency and/or asymptotic return than baselines? (Usually)
  • Q2: Can we verify that the two key ideas of POAM—agent modelling and use of data from uncontrolled agents—work as desired and contribute positively towards POAM’s performance? (Yes)
  • Q3: Does agent modelling improve generalization to previously unseen teammates? (Yes)
Full implementation details, including all hyperparameter values, along with additional empirical results, appear in Appendix. Our code is included with the submission and will be open-sourced upon publication.
@@ 6.1 Experimental Design
This subsection describes the experimental design, including the particular NAHT problem setting, training procedure, experimental domain, and baselines.
@@ A Practical Instantiation of NAHT
To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN [39], QMIX [32], IQL [40], IPPO, and MAPPO [43]. We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms are not compatible (Appendix A.4). This emergence of diverse teammate behaviors from training agents using different MARL algorithms under different seeds aligns with the findings from Strouse et al. [38].
Let $U_{train}$ denote the set of uncontrolled teammates used to train POAM and the baseline methods. $U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively. $U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training. The experimental results reported in 6.2 and 6.3 are computed with respect to $U_{train}$ only, while the experimental results in 6.4 use $U_{test}$.
@@
  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is use AHT algorithms that train only a single controlled agent, and copy these policies as many times as needed in the NAHT setting. To evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM that is trained identically to POAM, but where the number of controlled agents $N=1$ during training.
@@ 6.2 Main Results
Figure 3: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return. Results are computed over three trials, while the shaded region represents the 95% confidence interval.
@@
This section addresses Q1—that is, whether POAM improves asymptotic return and/or sample efficiency, compared to baselines. Figure 3 shows the learning curves of POAM and IPPO-NAHT, and the test returns achieved by the best naive MARL baseline and POAM-AHT, on all tasks.  [^4] See the Appendix for tables providing the performances of all naive MARL methods, across all tasks. All learning curves consist of the mean test returns across three trials, while the shaded regions reflect the 95% confidence intervals.
@@
We find that for all tasks, POAM outperforms baselines in terms of asymptotic return for three out of five tasks (pp, 5v6, 3s5z). For all tasks, POAM’s initial sample efficiency is similar to that of IPPO-NAHT for the first few million steps of training, after which POAM displays higher return. We attribute this to the initial costs incurred by learning team encoding vectors, which once learned, improves learning efficiency. IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method. Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to approximately the same returns on two out of five tasks (8v9 and 10v11). While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation. Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
Figure 4: Evolution of the within-episode mean squared error (left) and the within-episode probability of actions of modelled teammates (right), computed by the POAM agent over the course of training on the mpe-pp task.
@@
Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks. Overall, POAM discovers the most consistently performant policies compared to baseline methods, in a sample-efficient manner. We conclude that (1) agent modeling improves learning efficiency, and (2) direct duplication of AHT-trained policies is less effective than methods that co-train agents for NAHT.

### v3

@@ Abstract
Current approaches to learning cooperative multi-agent behaviors assume relatively restrictive settings. In standard fully cooperative multi-agent reinforcement learning, the learning algorithm controls all agents in the scenario, while in ad hoc teamwork, the learning algorithm usually assumes control over only a single agent in the scenario. However, many cooperative settings in the real world are much less restrictive. For example, in an autonomous driving scenario, a company might train its cars with the same learning algorithm, yet once on the road, these cars must cooperate with cars from another company. Towards expanding the class of scenarios that cooperative learning methods may optimally address, we introduce $N$-agent ad hoc teamwork (NAHT), where a set of autonomous agents must interact and cooperate with dynamically varying numbers and types of teammates. This paper formalizes the problem, and proposes the Policy Optimization with Agent Modelling (POAM) algorithm. POAM is a policy gradient, multi-agent reinforcement learning approach to the NAHT problem that enables adaptation to diverse teammate behaviors by learning representations of teammate behaviors. Empirical evaluation on tasks from the multi-agent particle environment and StarCraft II shows that POAM improves cooperative task returns compared to baseline approaches, and enables out-of-distribution generalization to unseen teammates.
@@ 3 NAHT Problem Formulation
Without loss of generality, let $C(\theta)=\{\pi_{i}^{\theta}(.|s)\}_{i=1}^{N}$ denote a set of $M$ controlled agent policies parameterized by $\theta$, such that a learning algorithm might optimize for $\theta$. Let the $\bm{\pi}^{(M)}$ indicate a team of $M$ agents and $\bm{\pi}^{(M)}\sim X(U,C)$ indicate sampling such a team from $U$ and $C$ via the team sampling procedure. The objective of the NAHT problem is to find parameters $\theta$, such that $C(\theta)$ maximizes the expected return in the presence of teammates from $U$:
@@ 4 The Need for Dedicated NAHT Algorithms
Define the following simple game for $M$ agents: at each turn, each agent $a_{i}$ picks one bit $b_{i}\in\{0,1\}$; at the end of each turn, all the bits are summed $s=\sum_{i}b_{i}$. The team wins if the sum of the chosen bits is exactly $1$. We denote the probability of winning by $P(s=1)$. Suppose the uncontrolled agents follow a policy that independently selects $1$ with probability $p=\frac{1}{M}$. [^3] Lemma A.1 shows that $p=\frac{1}{M}$ is the optimal value of $p$ for a team composed of such agents. In the following, we consider the three agent case, $M=3$, for simplicity.
In the AHT problem setting, a learning algorithm assumes control of only a single agent. Let $p_{\text{aht}}$ denote the probability with which the AHT agent selects 1. Given the aforementioned team of uncontrolled agents, we show that any value of $p_{\text{aht}}$ results in the same probability of winning, which occurs because the probability of winning, $P(s=1)=\frac{4}{9}$, is independent of $p_{\text{aht}}$ (Lemma A.2).
@@
Finally, we show there exists an NAHT policy that controls both agents and obtains a higher winning probability. Consider the policy where one controlled agent always plays 0, while the other plays 1 with probability $p_{\text{naht}}$. Lemma A.4 shows that the optimal $p_{\text{naht}}=1$, and the probability of winning $P(s=1)=\frac{2}{3}>\frac{4}{9}$. Thus, we have exhibited an NAHT scenario where an AHT policy that is optimal when $N=1$, performs worse than a simple NAHT joint policy in the $N=2$ setting. Empirical validation of the prior results are provided in Appendix A.5.1.
@@ 5 Policy Optimization with Agent Modeling (POAM)
Figure 2: POAM trains a single policy network $\pi^{\theta_{p}}$, which characterizes the behavior of all controlled agents (green), while uncontrolled agents (yellow) are drawn from $U$. Data from both controlled and uncontrolled agents is used to train the value network, $V^{\theta_{c}}_{i}$ while the policy is trained on data from the controlled agents only. The policy and value function are both conditioned on a learned team embedding vector, $e^{t}_{i}$.
This section describes the proposed Policy Optimization with Agent Modeling (POAM) method, which trains a collection of NAHT agents that can adaptively deal with different collections of unknown teammates. POAM relies on an agent modeling network to initially build an embedding vector characterizing teammates encountered during an interaction. Adaptive agent policies that can maximize the controlled agents’ returns are then learned by training a policy conditioned on the environment observation and team embedding vector. To enable controlling a varying number of agents while learning in a sample-efficient manner, POAM adopts the independent learning framework with full parameter sharing. The training processes for agent modeling and policy networks are described in Sections 5.1 and 5.2 respectively, while an illustration of how POAM trains NAHT agents is provided in Figure 2.
@@ 5.1 Agent Modeling Network
Designing adaptive policies that enable NAHT agents to achieve optimal returns against any team of uncontrolled agents drawn from some set $U$, requires information on the encountered team’s unknown behavior. However, in the absence of prior knowledge about uncontrolled teammates’ policies, local observations from a single timestep may not contain sufficient information regarding the encountered team. To circumvent this lack of information, POAM’s agent modeling network plays a crucial role in providing team embedding vectors that characterize the observed behavior of teammates in the encountered team.
@@
We identify two main criteria for desirable team embedding vectors. First, team embedding vectors should identify information regarding the unknown state and behavior of other agents in the environment (both controlled and uncontrolled). Second, team embedding vectors should ideally be computable from the sequence of local observations and actions of the team. Fulfilling both requirements provides an agent with useful information for decision-making in NAHT problems, even under partial observability.
@@
For each controlled agent, POAM produces informative team embedding vectors by training a model with an encoder-decoder architecture, illustrated by red components in Figure 2. For ease of presentation, the encoder-decoder models for controlled agent $i$ will be referred to without the index $i$. The encoder, $f^{\text{enc}}_{\theta^{e}}:H_{i}\mapsto\mathbb{R}^{n}$, is parameterized by $\theta^{e}$ and processes the modeling agent’s history of local observations and actions up to timestep $t$, $h_{i}^{t}=\{o_{i}^{k},a_{i}^{k-1}\}_{k=1}^{t}$, to compute a team embedding vector of dimension $n$, $e^{t}_{i}\in\mathbb{R}^{n}$ that characterizes the modeled agents. This reliance on local observations helps ensure that the agent modeling network can operate without having access to the environment state that is unavailable under partial observability. The team embedding vector is decoded by two decoder networks: the observation decoder, $f^{\text{dec}}_{\theta^{o}}:\mathbb{R}^{n}\mapsto O_{-i}$ and the action decoder, $f^{\text{dec}}_{\theta^{a}}:\mathbb{R}^{n}\mapsto\Delta(A_{-i})$. The decoder networks are respectively trained to predict the observations and actions of all other agents on the team at timestep $t$, $(o_{-i}^{t},a_{-i}^{t})$, to encourage $e^{t}_{i}$ to contain relevant information for the current NAHT agent’s decision-making process. While the observation decoder directly predicts the observed $-i$ observations, the action decoder predicts the parameters of a probability distribution over the $-i$ agents’ actions, $p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t})))$, where an appropriate distribution for $p$ should be selected by the system designer.
Concretely, agent $i$’s encoder-decoder model is trained to minimize a maximum likelihood loss over all teammates’ observations and actions, given its own local observations and actions. As the experimental setting in this paper considers continuous observations and discrete actions, the observation loss is a mean squared error loss, while the action loss is the negative log likelihood of the $-i$ agents’ actions, under the Categorical distribution.
$L_{\theta^{e},\theta^{o},\theta^{a}}(h_{i}^{t},o_{-i}^{t},a_{-i}^{t})=\Big{(}||f^{\text{dec}}_{\theta^{o}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))-o_{-i}^{t}||^{2}-\text{log}(p(a^{t}_{-i};f^{\text{dec}}_{\theta^{a}}(f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))))\Big{)}.$ (2)
@@ 5.2 Policy and Value Networks
The policy network of agent $i$, $\pi_{i}^{\theta^{p}}:H_{i}\times\mathbb{R}^{n}\mapsto\Delta(A_{i})$, is parameterized by $\theta^{p}$, and uses the NAHT agent’s local observation, $o^{t}_{i}$, and the team embedding from the encoder network, $e^{t}_{i}$, to compute a policy followed by the NAHT agents. Conditioning the policy network on $e^{t}_{i}$ allows an NAHT agent to change its behaviour based on the inferred characteristics of encountered agents. When training the policy network, we also rely on a value (or critic) network, $V_{i}^{\theta^{c}}:H_{i}\times\mathbb{R}^{n}\mapsto\mathbb{R}$, parameterized by $\theta^{c}$, which measures the expected returns given $h_{i}^{t}$, and $e^{t}_{i}$. The value network serves as a baseline to reduce the variance of the gradient updates, while conditioning on the learned teammate embeddings for similar reasons to the policy.
POAM then trains the policy and value networks using an approach based on the Independent PPO algorithm [45] (IPPO). IPPO is selected as the base MARL algorithm for two reasons. First, using an independent MARL method circumvents the need to deal with the changing number of agents resulting from environment openness. Second, IPPO has been demonstrated to be effective on various MARL tasks. To improve learning efficiency and enable information sharing between agents, full parameter sharing is employed for all neural networks. POAM trains the value network to produce accurate state value estimates by minimizing the following loss function:
$L_{\theta^{c}}(h_{i}^{t})=\dfrac{1}{2}\Big{(}V_{i}^{\theta^{c}}(h_{i}^{t},f^{\text{enc}}_{\theta^{e}}(h_{i}^{t}))-\hat{V_{i}}^{t}\Big{)}^{2},$ (3)
where $\hat{V_{i}}^{t}$ is the TD($\lambda$) return. The policy network is analogously trained to minimize the PPO loss function [35], but where the policy additionally conditions on the team embeddings.
@@
In the NAHT setting, we assume access to the joint observations and actions generated by the current team deployed in the environment at training time only, where the team consists of a mix of controlled and uncontrolled agents. This assumption provides an opportunity to learn useful cooperative behaviors more quickly, by bootstrapping based on transitions from the initially more competent, uncontrolled teammate policies, as also observed by Rahman et al. [31].
POAM leverages data from both controlled and uncontrolled agents to train the value network—in effect, treating the uncontrolled agents as exploration policies. Note that this aspect of POAM is a significant departure from PPO, which is a fully on-policy algorithm. Since the policy update is highly sensitive to off-policy data, only data from the controlled agents is used to train the policy network.
@@ 6 Experiments and Results
  • Q1: Does POAM learn to cope with uncontrolled teammates with higher sample efficiency and asymptotic return than baselines? (Usually)
  • Q2: Does POAM improve generalization to previously unseen and out-of-distribution teammates, compared to baselines? (Yes)
  • Q3: Can we verify that the two key ideas of POAM—agent modelling and use of data from uncontrolled agents—work as desired and contribute positively towards POAM’s performance? (Yes)
Full implementation details, including hyperparameter values and additional empirical results, appear in the Appendix. Our code is available at https://github.com/carolinewang01/naht.
@@ 6.1 Experimental Design
In the following, we summarize the experimental design, including the particular NAHT problem instance, training procedure, experimental domain, and baselines. Details on evaluation metrics are provided in Appendix A.3.3.
@@ A Practical Instantiation of NAHT
To generate a set of uncontrolled teammates $U$, the following MARL algorithms are used to train agent teams: VDN [41], QMIX [33], IQL [42], IPPO, and MAPPO [45]. We verify that the generated team behaviors are diverse by checking that (1) teams trained by the same algorithm learn non-compatible coordination conventions, and (2) teams trained by different algorithms also learn non-compatible coordination conventions (Appendix A.5.2). The emergence of diverse teammate behaviors from training agents using different MARL algorithms under different seeds aligns with the findings from Strouse et al. [40].
Let $U_{train}$ denote the set of uncontrolled teammates used to train all (N)AHT methods. $U_{train}$ consists of five teams, where each individual team is trained via VDN, QMIX, IQL, IPPO and MAPPO, respectively. $U_{test}$ consists of a set of holdout teams trained via the same MARL algorithms, but that have not been seen during training. The experimental results reported in Sections 6.2 and 6.4 are computed with respect to $U_{train}$ only, while the experimental results in Section 6.3 use $U_{test}$.
@@
  • POAM in the AHT setting (POAM-AHT): As considered in Section 4, a natural baseline approach to the NAHT problem is to use AHT algorithms that train only a single controlled agent, and copy these policies as many times as needed in the NAHT setting. To evaluate the intuition that AHT policies do not suffice for the NAHT problem setting, we consider an AHT version of POAM that is trained identically to POAM, but where the number of controlled agents is always one ($N=1$) during training. Note that POAM-AHT is equivalent to the AHT algorithm introduced by Papoudakis et al. [29], LIAM.
@@ 6.2 Main Results
Figure 3: POAM consistently improves over the baselines of IPPO-NAHT, POAM-AHT, and the best naive MARL baseline in all tasks, in either sample efficiency or asymptotic return.
@@
This section addresses Q1—that is, whether POAM learns to cope with uncontrolled teammates with greater sample efficiency or asymptotic returns, compared to baselines. Figure 3 shows the learning curves of POAM and IPPO-NAHT, and the test returns achieved by the best naive MARL baseline and POAM-AHT, on all tasks. [^4] See the Appendix for tables providing the performances of all naive MARL methods, across all tasks. All learning curves consist of the mean test returns across five trials, while the shaded regions reflect the 95% confidence intervals.
@@
We find that for all tasks, POAM outperforms baselines in terms of asymptotic return for three out of five tasks (mpe-pp, 5v6, 3s5z). For all tasks, POAM’s initial sample efficiency is similar to that of IPPO-NAHT for the first few million steps of training, after which POAM displays higher return. We attribute the initial similarity in sample efficiency to the initial cost incurred by learning team embedding vectors, which once learned, improves learning efficiency. IPPO-NAHT, which can be viewed as an ablation of POAM with no agent modeling, is the next best performing method. Although IPPO-NAHT has generally poorer sample efficiency than POAM, the method converges to approximately the same returns on two out of five tasks (8v9 and 10v11). While the agent modeling module of POAM provides team embedding vectors to the policy learning process, the embeddings are themselves produced from each agent’s own observation. Since no additional information is provided to POAM agents, it is unsurprising that IPPO-NAHT can converge to similar solutions as POAM, given enough training steps.
Figure 4: Test returns achieved by POAM and IPPO-NAHT, when paired with out-of-distribution teammates. POAM has improved generalization to OOD teammates, compared with IPPO-NAHT.
@@
Finally, while the best naive MARL baseline and POAM-AHT learn good solutions on some tasks, neither method consistently performs well across all tasks. Overall, POAM discovers the most consistently performant policies compared to baseline methods, in a relatively sample-efficient manner. We conclude that (1) agent modeling improves learning efficiency, and (2) direct duplication of AHT-trained policies is less effective than methods that co-train agents for NAHT.
6.3 Out of Distribution Generalization
The AHT literature commonly assumes that uncontrolled teammates of interest may be interacted directly with during training [29, 6, 31]; experiments in the prior section were conducted under this assumption. However, in realistic scenarios, it may be challenging to enumerate all teammates likely to be encountered in the wild. This section examines the effectiveness of POAM under a true NAHT evaluation scenario, where POAM agents must coordinate with teammates that were not available at training time and are out-of-distribution (OOD) (Q2). Here, OOD teammates are created by running MARL algorithms with different random seeds than those used to generate train-time teammates.
Figures 4 and 10 show the mean and 95% confidence intervals of the test return achieved by POAM, compared to IPPO-NAHT, when the algorithm in question is paired with previously unseen seeds of IPPO, IQL, MAPPO, QMIX, and VDN. For each type of teammate, the performance of IPPO-NAHT/POAM against the exact teammates seen during training is shown as the in-distribution baseline. Both POAM and IPPO-NAHT consistently exhibit reduced performance against the OOD teammates, compared to their respective in-distribution performances. In three out of five tasks (mpe, 5v6, 3s5z), POAM has a significantly higher return than IPPO-NAHT, while the remaining two tasks exhibit a smaller improvement. In Appendix A.5.3, similar findings are presented with an alternative OOD teammate generation strategy, where the set of five MARL algorithms used to generate uncontrolled teammates (IPPO, IQL, MAPPO, QMIX, VDN) are divided into train/test sets.